OpenStudy (anonymous):

True or False? If tan(θ) = 3/5, then sin(θ) = 3 and cos(θ) = 5.

OpenStudy (anonymous):

true; since \[\tan(\theta)=\sin(\theta)/\cos(\theta)\]

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

I got it wrong, but I think that it was because since it would be a 3,4,5 triangle (it has to be a right triangle), then sin(θ) would have to be 3/4 and cos(θ) would have to be 5/4. It is the only explanation I have...

OpenStudy (midhun.madhu1987):